OpenStudy (anonymous):

Part 1: Use the quadratic formula to solve x2 + 8x = –2. (3 points) Part 2: Using complete sentences, explain the process you used. (2 points)

OpenStudy (anonymous):

The quadratic formula is: x1 = [-b+sqrt(b^2 - 4ac)]/2a x2= [-b-sqrt(b^2 - 4ac)]/2a The quadratic is x^2 + 8x + 2 = 0 a,b,c are the coefficients of quadratic a = 1, b = 8 and c = 2 x1 = [-8+sqrt(64 - 8)]/2 x1 = (-8+sqrt56)/2 x1 = (-8+2sqrt(14))/2 x1 = -4+sqrt14 x2 = -4-sqrt14

OpenStudy (anonymous):

Which u will now have to do yourself so that u can explain the process u used...